# PACS Collaboration

### v1

@@ III.2 Pseudoscalar form factor $\widetilde{G}_{P}(q^{2})$
The $\widetilde{G}_{P}$ form factor is extracted from Eq. (30) as a function of the current insertion time $t$. As shown in Fig.7, similar to the $F_{P}$ form factor, the leading $\pi N$ subtraction method eliminates the slight convex shape associated with the excited-state contribution and yields a plateau behavior consistent with the PPD model for each $t_{\mathrm{sep}}$. As in the case of $F_{P}(q^{2})$, the values of $\widetilde{G}_{P}(q^{2})$ are easily determined in the standard manner using the uncorrelated constant fit. Figure 8 shows that there is no $t_{\mathrm{sep}}$ dependence of $\widetilde{G}_{P}$ for all $q^{2}$ in either ensemble.
@@
Figure 7: Using the leading $\pi N$ subtraction method as defined in Eq. (30), the values of $\widetilde{G}_{P}$ calculated from PACS10/L160 with $t_{\mathrm{sep}}/a=13$ (diamonds), $16$ (squares) and $19$ (circles) for seven non-zero momentum transfers (labeled from Q1 to Q7) as functions of the current insertion time slice $t$. The horizontal dashed line together with yellow band in each panel is calculated from the PPD model ($\widetilde{G}_{P}^{\rm PPD}(q^{2})$) given in Eq. (5).
@@ III.3 Induced pseudoscalar coupling $g_{P}^{*}$ and the pion-nucleon coupling $g_{\pi NN}$
In this study, we confirm that the value of $k_{\mathrm{max}}=3$ for the $128^{4}$ and $160^{4}$ lattices ($k_{\mathrm{max}}=4$ for the $64^{4}$ lattice) is large enough to guarantee that the $z$-expansion analysis makes a model-independent fit. We then extrapolate $F_{P}(q^{2})$ at $q^{2}=q_{0}^{2}$ to evaluate the value of $g_{P}^{\ast}$. The results of $g_{P}^{\ast}$ obtained from the coarse ($128^{4}$) and fine ($160^{4}$) lattices are summarized in Table 4. We also determine the value of $g_{P}^{\ast}$ using the smaller $64^{4}$ lattice ensemble at the coarse lattice spacing.
@@
$\beta$ Ensemble $t_{\mathrm{sep}}/a$ of data $g_{P}^{\ast}$ $g_{\pi NN}$
@@
Since the accessible $q^{2}_{\mathrm{min}}$ moves away from $q^{2}=q_{0}^{2}$ or $q^{2}=0$ as the spatial volume decreases, the convergence of the $z$-expansion for the $64^{4}$ lattice results was expected to be much worse than for the larger volume calculation. However, if $k_{\mathrm{max}}=4$ is chosen, the convergence of the $z$-expansion is well controlled. As a result, the obtained values of $g_{P}^{\ast}$ agree within their statistical error with those of the $128^{4}$ lattice ensemble as listed in Table 4.
@@
Next, we would like to evaluate another quantity obtained from the $F_{P}$ form factor. From the phenomenological point of view, the residue of the pion pole in the $F_{P}$ form factor is related to the pion nucleon coupling $g_{\pi NN}$ Nambu:1960xd . The $F_{P}$ form factor should be expressed as
@@
Figure 11 shows the $t_{\mathrm{sep}}$ dependence of the evaluated values of the (renormalized) pion-nucleon coupling $g_{\pi NN}$. As is in the case with $g_{A}$ and $g_{P}^{\ast}$, the systematic uncertainties due to excited-state contamination are well controlled, while both finite volume and lattice discretization effects do not manifest beyond statistical uncertainties. The horizontal dashed line together with gray band indicates the iso-spin average value of the experimental results Babenko:2016idp ; Limkaisang:2001yz as $g_{\pi NN}^{2}=\frac{1}{3}\left(g_{\pi^{0}NN}^{2}+2g_{\pi^{\pm}NN}^{2}\right)$. Our results for $g_{\pi NN}$ again agree well with the experimental results Babenko:2016idp ; Limkaisang:2001yz , with good accuracy.
@@
Similar to the case of $g_{A}$ cited in Refs. Tsuji:2023llh ; Tsuji:2022ric , we quote our best estimates of $g_{P}^{\ast}$ and $g_{\pi NN}$ under the condition of $t_{{\mathrm{sep}}}\gtrsim 1.2\ \mathrm{fm}$. We then evaluate the combined values of $g_{P}^{\ast}$ and $g_{\pi NN}$ with $t_{\mathrm{sep}}/a=\{14,16\}$ from both $128^{4}$ and $64^{4}$ calculations for the coarse lattice result, while the results obtained with a single data set of $t_{\mathrm{sep}}/a=19$ are chosen for the fine lattice as our best estimates. Our final results for $g_{A}$, $g_{P}^{\ast}$ and $g_{\pi NN}$ are summarized in Table 6 together with the measured nucleon masses. The first error is a statistical one, while the others are systematic ones. The second error is a systematic error stemming from the error of $Z_{A}^{\mathrm{SF}}$ for the renormalization. The third error is evaluated by the difference between two analyses using either the $t_{{\mathrm{sep}}}\gtrsim 1.2\ \mathrm{fm}$ or $t_{{\mathrm{sep}}}\gtrsim 1.0\ \mathrm{fm}$ condition for the selection of data sets in a combined analysis. Two systematic errors are smaller than the statistical errors of either quantity at two lattice spacings. Recall that the systematic error due to contamination from exited-states is fairly well under controlled.
Figure 10: The source-sink separation ($t_{\mathrm{sep}}$) dependence of the renormalized values of the induced pseudoscalar coupling $g_{P}^{*}$. The horizontal axis gives $t_{\mathrm{sep}}$ in physical units. The diamond and circle symbols are results for the $128^{4}$ (coarse) and $160^{4}$ (fine) lattices. The horizontal dashed line together with gray band denotes the experimental result MuCap:2007tkq ; MuCap:2012lei .
Figure 11: Same as Fig. 10 for the pion-nucleon coupling $g_{\pi NN}$. The iso-spin average is used for the experimental result as $g^{2}_{\pi NN}=\frac{1}{3}\left(g_{\pi^{0}NN}^{2}+2g_{\pi^{\pm}NN}^{2}\right)$ Babenko:2016idp ; Limkaisang:2001yz .
@@
We finally compare our results of the renormalized value of $g_{P}^{\ast}$ (left) and $g_{\pi NN}$ (right) together with results form the recent lattice QCD simulations RQCD:2019jai ; Gupta:2024qip ; Park:2021ypf ; Jang:2023zts ; Alexandrou:2023qbg and experimental results MuCap:2007tkq ; MuCap:2012lei ; Babenko:2016idp ; Limkaisang:2001yz in Fig. 12. The inner error bars represent the statistical uncertainties, while the outer ones represent the total uncertainties given by adding the statistical and systematic errors in quadrature. Our results show that the systematic errors are so well controlled that the inner and outer errors are not visible in Fig 12.
It is important to note that our results obtained at two lattice spacings are derived solely from the physical point simulations, which would suffer from large statistical fluctuations. In contrast, the other lattice results except for the ETMC 2023 result Alexandrou:2023qbg are given by the combined data that incorporates the data obtained from simulations at the heavier pion masses RQCD:2019jai ; Park:2021ypf ; Jang:2023zts . Nevertheless, our results obtained at two lattice spacings are significantly smaller statistical and systematic errors for both $g_{P}^{\ast}$ and $g_{\pi NN}$ compared to the other lattice results RQCD:2019jai ; Gupta:2024qip ; Park:2021ypf ; Jang:2023zts ; Alexandrou:2023qbg . Our subtraction method can get more accurate results of two target quantities, $g_{P}^{\ast}$ and $g_{\pi NN}$, comparing the multi-state fit approach that was used in other groups. See, Appendix D for more details.
@@ IV Summary
In this study, we thus propose a simple subtraction method for removing the so-called leading $\pi N$-state contamination induced by the pion-pole structure appears in $F_{P}(q^{2})$ and $G_{P}(q^{2})$. The leading $\pi N$ state contribution in the three-point function of the spatial axial-vector current $A_{i}$ was found to be eliminated by a specific linear combination of the two time derivatives of the correlator ratios composed of the $A_{4}$ and $A_{i}$ currents. The leading $\pi N$ subtraction method can be easily applied to existing data without the need for new calculations, provided that the excited-state contribution, which appears in the standard ratio method, is sufficiently hindered for $F_{A}(q^{2})$ by applying the appropriate smearing to the nucleon operator.
@@
The results of both $F_{P}(q^{2})$ and $G_{P}(q^{2})$ obtained from the new analysis surprisingly agree well with the prediction of the PPD model in the range of $0.01\lesssim q^{2}\lesssim 0.12$ $[{\rm GeV}^{2}]$ measured in this paper. There is experimental information coming from two experiments for $F_{P}(q^{2})$. One is the experiment of the muon capture by the proton, the other is the pion electro-production experiment. Our results agree well with these experimental results as well.
@@
$g_{P}^{\ast}$ =$=\left\{\begin{array}[]{lr}\makebox[80.00012pt][l]{8.53(23)(6)(4)}&\beta=1.82\\ \makebox[80.00012pt][l]{8.74(16)(2)(13)}&\beta=2.00\\ \end{array}\right.$ (38)
$g_{\pi NN}$ =$=\left\{\begin{array}[]{lcr}\makebox[80.00012pt][l]{12.80(41)(9)(4)}&\beta=1.82\cr\makebox[80.00012pt][l]{13.24(28)(3)(14)}&\beta=2.00\cr\end{array}\right.$ (41)
@@
Since our lattice simulations are currently running at the third lattice spacing to perform the continuum limit extrapolation Tsuji:2024scy , we have estimated the systematic uncertainties associated with the finite lattice spacing for $g_{P}^{\ast}$ and $g_{\pi NN}$ from the difference between from the current results obtained at two lattice spacings. It was found that the finite lattice spacing effects on $g_{P}^{\ast}$ and $g_{\pi NN}$ are kept below the statistical error of less than 4%, while all results are consistent with respective experimental values. Therefore, the lattice discretization effects on $g_{P}^{\ast}$ and $g_{\pi NN}$ are negligibly small in our calculations, similar to the case of $g_{A}$ as reported in Ref Tsuji:2023llh .
A comprehensive study that employs the current results for $F_{P}(q^{2})$ and $G_{P}(q^{2})$ along with $F_{A}(q^{2})$ to verify the generalized GT relation and the PPD model is outlined in a separate paper NewPaper and not included in this paper. Although the axial radius $r_{A}$ is most relevant for current and future neutrino experiments, there is a possible discretization error of about 10% in this quantity as pointed out in Ref. Tsuji:2023llh . Therefore, it is necessary to continue our lattice QCD calculations for the axial structure of the nucleon with the third lattice spacing toward the continuum limit. Such a study is currently underway Tsuji:2024scy .
@@ Acknowledgements.
We would like to thank members of the PACS collaboration for useful discussions. K.-I. I. is supported in part by MEXT as “Feasibility studies for the next-generation computing infrastructure”. K. S. is supported by JST, The Establishment of University Fellowships towards the creation of Science Technology Innovation, Grant Number JPMJFS2106. We also thank Y. Namekawa for his careful reading of the manuscript. Numerical calculations in this work were performed on Oakforest-PACS in Joint Center for Advanced High Performance Computing (JCAHPC) and Cygnus and Pegasus in Center for Computational Sciences at University of Tsukuba under Multidisciplinary Cooperative Research Program of Center for Computational Sciences, University of Tsukuba, and Wisteria/BDEC-01 in the Information Technology Center, the University of Tokyo. This research also used computational resources of the K computer (Project ID: hp180126) and the Supercomputer Fugaku (Project ID: hp20018, hp210088, hp230007, hp230199, hp240028, hp240207, hp250037) provided by RIKEN Center for Computational Science (R-CCS), as well as Oakforest-PACS (Project ID: hp170022, hp180051, hp180072, hp190025, hp190081, hp200062), Wisteria/BDEC-01 Odyssey (Project ID: hp220050) provided by the Information Technology Center of the University of Tokyo / JCAHPC. The calculation employed OpenQCD system(http://luscher.web.cern.ch/luscher/openQCD/). This work is supported by the JLDG constructed over the SINET6 of NII, This work was also supported in part by Grants-in-Aid for Scientific Research from the Ministry of Education, Culture, Sports, Science and Technology (Nos. 22K03612, 23H01195, 23K03428, 25KJ0404) and MEXT as “Program for Promoting Researches on the Supercomputer Fugaku” (Search for physics beyond the standard model using large-scale lattice QCD simulation and development of AI technology toward next-generation lattice QCD; Grant Number JPMXP1020230409).

### v2

@@ III.2 Pseudoscalar form factor $\widetilde{G}_{P}(q^{2})$
The $\widetilde{G}_{P}$ form factor is extracted from Eq. (30) as a function of the current insertion time $t$. As shown in Fig.7, similar to the $F_{P}$ form factor, the leading $\pi N$ subtraction method eliminates the slight convex shape associated with the excited-state contribution and yields a plateau behavior consistent with the PPD model for each $t_{\mathrm{sep}}$. As in the case of $F_{P}(q^{2})$, the values of $\widetilde{G}_{P}(q^{2})$ are easily determined in the standard manner using the uncorrelated constant fit. Figure 8 shows that there is no $t_{\mathrm{sep}}$ dependence of $\widetilde{G}_{P}$ for all $q^{2}$ in either ensemble. Furthermore, there is good agreement with the results of the summation method.
@@
Figure 7: Using the leading $\pi N$ subtraction method as defined in Eq. (30), the values of $\widetilde{G}_{P}$ calculated from PACS10/L160 with $t_{\mathrm{sep}}/a=13$ (diamonds), $16$ (squares) and $19$ (circles) are shown for seven nonzero momentum transfers (labeled from Q1 to Q7) as functions of the current insertion time slice $t$. The horizontal dashed line together with yellow band in each panel is calculated from the PPD model ($\widetilde{G}_{P}^{\rm PPD}(q^{2})$) given in Eq. (5).
@@ III.3 Induced pseudoscalar coupling $g_{P}^{*}$ and the pion-nucleon coupling $g_{\pi NN}$
In this study, we confirm that the value of $k_{\mathrm{max}}=3$ for the $128^{4}$ and $160^{4}$ lattices ($k_{\mathrm{max}}=4$ for the $64^{4}$ lattice) is large enough to guarantee that the $z$-expansion analysis makes a model-independent fit.  [^4] The model independence is verified by the convergence of the $z$-expansion under the variation of $k_{\rm max}$ for $G(q^{2})$. This procedure is analogous to the one described in Ref. Tsuji:2023llh . We then extrapolate $F_{P}(q^{2})$ at $q^{2}=q_{0}^{2}$ to evaluate the value of $g_{P}^{\ast}$. The results of $g_{P}^{\ast}$ obtained from the coarse ($128^{4}$) and fine ($160^{4}$) lattices are summarized in Table 4. We also determine the value of $g_{P}^{\ast}$ using the smaller $64^{4}$ lattice ensemble at the coarse lattice spacing.
@@
$\beta$ Ensemble $t_{\mathrm{sep}}/a$ # of data $g_{P}^{\ast}$ $g_{\pi NN}$
@@
Since the accessible $q^{2}_{\mathrm{min}}$ moves away from $q^{2}=q_{0}^{2}$ or $q^{2}=0$ as the spatial volume decreases, the convergence of the $z$-expansion for the $64^{4}$ lattice results was expected to be much worse than for the larger volume calculation. However, if $k_{\mathrm{max}}=4$ is chosen, the convergence of the $z$-expansion is well controlled. As a result, the obtained values of $g_{P}^{\ast}$ agree within their statistical errors with those of the $128^{4}$ lattice ensemble as listed in Table 4.
@@
Next, we would like to evaluate another quantity obtained from the $F_{P}$ form factor. From the phenomenological point of view, the residue of the pion pole in the $F_{P}$ form factor is related to the pion-nucleon coupling $g_{\pi NN}$ Nambu:1960xd . The $F_{P}$ form factor should be expressed as
@@
Figure 11 shows the $t_{\mathrm{sep}}$ dependence of the evaluated values of the (renormalized) pion-nucleon coupling $g_{\pi NN}$. As is in the case with $g_{A}$ and $g_{P}^{\ast}$, the systematic uncertainties due to excited-state contamination are well controlled, while both finite volume and lattice discretization effects do not manifest beyond statistical uncertainties. The horizontal dashed line together with gray band indicates the isospin average value of the experimental results Babenko:2016idp ; Limkaisang:2001yz ; Reinert:2020mcu as $g_{\pi NN}^{2}=\frac{1}{3}\left(g_{\pi^{0}NN}^{2}+2g_{\pi^{\pm}NN}^{2}\right)$. Our results for $g_{\pi NN}$ again agree well with the experimental results Babenko:2016idp ; Limkaisang:2001yz ; Reinert:2020mcu , with good accuracy.
@@
Similar to the case of $g_{A}$ cited in Refs. Tsuji:2023llh ; Tsuji:2022ric , we quote our best estimates of $g_{P}^{\ast}$ and $g_{\pi NN}$ under the condition of $t_{{\mathrm{sep}}}\gtrsim 1.2\ \mathrm{fm}$. We then evaluate the combined values of $g_{P}^{\ast}$ and $g_{\pi NN}$ with $t_{\mathrm{sep}}/a=\{14,16\}$ from both $128^{4}$ and $64^{4}$ calculations for the coarse lattice result, while the results obtained with a single data set of $t_{\mathrm{sep}}/a=19$ are chosen for the fine lattice as our best estimates. Our final results for $g_{A}$, $g_{P}^{\ast}$ and $g_{\pi NN}$ are summarized in Table 6 together with the measured nucleon masses. The first error is a statistical one, while the others are systematic ones. The second error is a systematic error stemming from the error of $Z_{A}^{\mathrm{SF}}$ for the renormalization. The third error is evaluated by the difference between two analyses using either the $t_{{\mathrm{sep}}}\gtrsim 1.2\ \mathrm{fm}$ or $t_{{\mathrm{sep}}}\gtrsim 1.0\ \mathrm{fm}$ condition for the selection of datasets in a combined analysis. Two systematic errors are smaller than the statistical errors of either quantity at two lattice spacings. Recall that the systematic error due to contamination from excited states is quite well controlled.
Figure 10: Source-sink separation ($t_{\mathrm{sep}}$) dependence of the renormalized values of the induced pseudoscalar coupling $g_{P}^{*}$. The horizontal axis gives $t_{\mathrm{sep}}$ in physical units. The diamond and circle symbols are results for the $128^{4}$ (coarse) and $160^{4}$ (fine) lattices. The horizontal dashed line together with gray band denotes the experimental result MuCap:2007tkq ; MuCap:2012lei .
Figure 11: Same as Fig. 10 for the pion-nucleon coupling $g_{\pi NN}$. The isospin average is used for the experimental result as $g^{2}_{\pi NN}=\frac{1}{3}\left(g_{\pi^{0}NN}^{2}+2g_{\pi^{\pm}NN}^{2}\right)$ Babenko:2016idp ; Limkaisang:2001yz .
@@
We finally compare our results of the renormalized value of $g_{P}^{\ast}$ (left) and $g_{\pi NN}$ (right) together with results form the recent lattice QCD simulations RQCD:2019jai ; Gupta:2024qip ; Park:2021ypf ; Jang:2023zts ; Alexandrou:2023qbg and experimental results MuCap:2007tkq ; MuCap:2012lei ; Babenko:2016idp ; Limkaisang:2001yz in Fig. 12. The inner error bars represent the statistical uncertainties, while the outer ones represent the total uncertainties given by adding the statistical and systematic errors in quadrature. Since our simulations are accomplished at only two lattice spacings, we have not yet performed the continuum limit extrapolation. Therefore, the two results obtained at different lattice spacings are plotted individually in Fig. 12. Our results show that the systematic errors are so well controlled that the inner and outer errors are not visible in Fig. 12. It is important to note that our two results are derived solely from the physical point simulations, which would suffer from large statistical fluctuations. In contrast, except for the Extended Twisted Mass Collaboration (ETMC) result Alexandrou:2023qbg , the other lattice results are given by the chiral-continuum extrapolation, which incorporates data obtained from simulations at heavier pion masses RQCD:2019jai ; Park:2021ypf ; Jang:2023zts .
@@ IV Summary
In this study, we thus propose a simple subtraction method for removing the so-called leading $\pi N$-state contamination induced by the pion-pole structure appears in $F_{P}(q^{2})$ and $G_{P}(q^{2})$. The leading $\pi N$-state contribution in the three-point function of the spatial axial-vector current $A_{i}$ was found to be eliminated by a specific linear combination of the two time derivatives of the correlator ratios composed of the $A_{4}$ and $A_{i}$ currents. The leading $\pi N$ subtraction method can be easily applied to existing data without the need for new calculations, provided that the excited-state contribution, which appears in the standard ratio method, is sufficiently hindered for $F_{A}(q^{2})$ by applying the appropriate smearing to the nucleon operator.
@@
The results of both $F_{P}(q^{2})$ and $G_{P}(q^{2})$ obtained from the new analysis surprisingly agree well with the prediction of the PPD model in the range of $0.01\lesssim q^{2}\lesssim 0.12$ ${\rm GeV}^{2}]$ measured in this paper. There is experimental information coming from two experiments for $F_{P}(q^{2})$. One is the experiment of the muon capture by the proton, the other is the pion electro-production experiment. Our results also agree well with these experimental results.
@@
[…]
$g_{\pi NN}$ $=\left\{\begin{array}[]{lcr}\makebox[80.00012pt][l]{12.80(41)(9)(4)}&\beta=1.82\cr\makebox[80.00012pt][l]{13.24(28)(3)(14)}&\beta=2.00\cr\end{array}\right$. (41)
@@
Since our lattice simulations are currently running at the third lattice spacing to perform the continuum limit extrapolation Tsuji:2024scy , we may estimate the size of the systematic uncertainties associated with the finite lattice spacing for $g_{P}^{\ast}$ and $g_{\pi NN}$ by comparing the current results obtained at two different lattice spacings. Similar to the case of $g_{A}$ as reported in Ref Tsuji:2023llh , this analysis reveals that the discrepancy between the two results obtained at different lattice spacings is smaller than their statistical errors for both values of $g_{P}^{\ast}$ and $g_{\pi NN}$. Furthermore, taking into account the preliminary results obtained at the third lattice spacing Tsuji:2024scy , we expect that the lattice discretization error will be maintained at a level of a few percent for both $g_{P}^{\ast}$ and $g_{\pi NN}$.
A comprehensive study that employs the current results for $F_{P}(q^{2})$ and $G_{P}(q^{2})$ along with $F_{A}(q^{2})$ to verify the generalized GT relation and the PPD model is outlined in a separate paper Tsuji:2025quu and not included in this paper. Although the axial radius $r_{A}$ is most relevant for current and future neutrino experiments, there is a possible discretization error of about 10% in this quantity as pointed out in Ref. Tsuji:2023llh . Therefore, it is necessary to continue our lattice QCD calculations for the axial structure of the nucleon with the third lattice spacing toward the continuum limit. Such a study is underway Tsuji:2024scy .
@@ Acknowledgements.
We would like to thank members of the PACS collaboration for useful discussions. K.-I. I. is supported in part by MEXT as “Feasibility studies for the next-generation computing infrastructure”. K. S. is supported by JST, The Establishment of University Fellowships towards the creation of Science Technology Innovation, Grant Number JPMJFS2106. We also thank Y. Namekawa for his careful reading of the manuscript. Numerical calculations in this work were performed on Oakforest-PACS in Joint Center for Advanced High Performance Computing (JCAHPC) and Cygnus and Pegasus in Center for Computational Sciences at University of Tsukuba under Multidisciplinary Cooperative Research Program of Center for Computational Sciences, University of Tsukuba, and Wisteria/BDEC-01 in the Information Technology Center, the University of Tokyo. This research also used computational resources of the K computer (Project ID: hp180126) and the Supercomputer Fugaku (Project ID: hp20018, hp210088, hp230007, hp230199, hp240028, hp240207, hp250037) provided by RIKEN Center for Computational Science (R-CCS), as well as Oakforest-PACS (Project ID: hp170022, hp180051, hp180072, hp190025, hp190081, hp200062), Wisteria/BDEC-01 Odyssey (Project ID: hp220050) provided by the Information Technology Center of the University of Tokyo/JCAHPC. The calculation employed OpenQCD system OpenQCD . This work is supported by the Japan Lattice Data Grid (JLDG) constructed over the SINET6 of NII. This work was also supported in part by Grants-in-Aid for Scientific Research from the Ministry of Education, Culture, Sports, Science and Technology (Nos. 22K03612, 23H01195, 23K03428, 25KJ0404) and MEXT as “Program for Promoting Researches on the Supercomputer Fugaku” (Search for physics beyond the standard model using large-scale lattice QCD simulation and development of AI technology toward next-generation lattice QCD; Grant Number JPMXP1020230409).